Of the 200 candidates who were interviewed for a position at a call center, 100 had a two-wheeler, 70 had a credit card and 140 had a mobile phone. 40 of them had both, a two-wheeler and a credit card, 30 had both, a credit card and a mobile phone and 60 had both, a two wheeler and mobile phone and 10 had all three. How many candidates had none of the three?

Grade:

4 Answers

raj kaushik
15 years ago
Number of candidates who had none of the three = Total number of candidates - number of candidates who had at least one of three devices.

Total number of candidates = 200.

Number of candidates who had at least one of the three = A U B U C, where A is the set of those who have a two wheeler, B the set of those who have a credit card and C the set of those who have a mobile phone.

We know that AUBUC = A + B + C - {A n B + B n C + C n A} + A n B n C

Therefore, AUBUC = 100 + 70 + 140 - {40 + 30 + 60} + 10

Or AUBUC = 190.

As 190 candidates who attended the interview had at least one of the three gadgets, 200 - 190 = 10 candidates had none of three.
raja kapoor
15 years ago

THANX

sakshi zutshi
11 years ago

A computer password consists of a letter of alphabet followed by three or four digits. Find the total number of passwords that can be formed.

anuj
7 years ago
Totally there are 200